OpenStudy (brucebaner):

A balloon’s circumference decreases at a constant rate. The graph shows the balloon’s circumference over time. What is the slope of the line that represents the situation? Enter the slope as a decimal.

OpenStudy (brucebaner):

@AriPotta

OpenStudy (aripotta):

take those two points (5,5) and (10,1) and plug them into the slope formula:\[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (brucebaner):

okay

OpenStudy (aripotta):

did you get it?

OpenStudy (brucebaner):

no

OpenStudy (aripotta):

ok so you have the points (5,5) and (10,1) so when you put them into the formula, you get \[\frac{ 5-1 }{ 5-10 }\]can you work it out to get the answer?

OpenStudy (brucebaner):

0.8

OpenStudy (aripotta):

it'll be negative tho

OpenStudy (brucebaner):

okay

OpenStudy (aripotta):

\[\frac{ 5-1 }{ 5-10}=~\frac{ 4 }{ -5 }\] that's why it's negative

OpenStudy (brucebaner):

so its -0.8 or -4/5

OpenStudy (aripotta):

yep. it says put it in as a decimal, so -0.8

OpenStudy (brucebaner):

thx

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\)